OpenStudy (anonymous):

Can someone quickly check 6 questions for me correct me if I did any wrong...I need to finish soon so I just want someone to correct me.

OpenStudy (the_fizicx99):

First one is wrong, I'm not sure on the second one. Keep forgetting radicals U.u

OpenStudy (anonymous):

Yeah Radicals was the hardest thing for me to learn

OpenStudy (the_fizicx99):

Suppose you're given 16 ^ 7/9 The numerator is the exponent of 16 the 9 is the root,

OpenStudy (the_fizicx99):

Decent diagram of what it should be.

OpenStudy (nikato):

2 is D.

OpenStudy (nikato):

If u rewrite it, it is 7^1/3 ------- 7^1/5 Using this rule, http://www.regentsprep.org/regents/math/algebra/ao5/rule_pdiv.htm U get 7^2/15

OpenStudy (anonymous):

is 1 is A? and thank you Nikato.

OpenStudy (nikato):

Yes 1 is A

OpenStudy (the_fizicx99):

Yeah 1 is A

OpenStudy (the_fizicx99):

I think you're right, 4 square is 2 and 25 squared is 5

OpenStudy (nikato):

Yea. Both look correct

OpenStudy (the_fizicx99):

It's decay because its below 1, anything below 1 is decay.. anything above 0.99 is growth

OpenStudy (anonymous):

Okay thank you

OpenStudy (the_fizicx99):

np

OpenStudy (the_fizicx99):

OHH WAIT I was confused!! See the y intercept of g(x) is when x is 0, when x is 0, the y intercept is left alone :D So because the y intercept of f(x) is 3 and the y intercept of g(x) is 6 then its 2 times the y intercept of f(x). 3 * 2 = 6

OpenStudy (the_fizicx99):

The y intercept of g(x) is 2 times the y intercept of f(x) ._.

OpenStudy (anonymous):

Wait so when there is a 0 the y intercept is left alone?

OpenStudy (the_fizicx99):

When x is 0, thats how you find the y intercept. To find the y intercept, set x for 0 To find the x intercept, set y for 0

OpenStudy (the_fizicx99):

Suppose your asked to find the y intercept of 3x + 2y = 12 Set x for 0, 3(0) + 2y = 12 /2 /2 y = 6 You can pretty much find the x from plugging in your y, 3x + 2(6) = 12 3x + 12 = 12 -12 -12 3x = 0 /3 /3 x = 0 3(0) + 2(6) = 12 0 + 12 = 12 12 = 12 :D

OpenStudy (anonymous):

Okay thank you this is very helpful! You're really good! I'm so sorry to bother you but i have to get a good grade on this mind checking like 2 more? so sorry!!

OpenStudy (the_fizicx99):

Sure, i gtg soon :p like next 10 min

OpenStudy (mertsj):

What did you put for number 5?

OpenStudy (anonymous):

C

OpenStudy (nikato):

C is correct. the first function is f(x)=5^2 second is g(x)=-3 and then together 5^2 -3

OpenStudy (nikato):

for the other one, change it to slope intercept form and grpah it so y=-3x+10. can u try grphing this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Is it C? sorry I might have graphed it wrong.. I am not really good at this.

OpenStudy (nikato):

no. its not right. however part of it is right. the graph does not contain the origin, however, it should be a slolid line when grapging, the >= sign means solid line where as just > means dotted line

OpenStudy (anonymous):

Ohh thank you!! (:

OpenStudy (nikato):

yea. and just to let u know, ur graph should have look something like this

OpenStudy (nikato):

https://www.desmos.com/calculator

OpenStudy (nikato):

oh nevermind, but if u want to look at it, just type in ur function

OpenStudy (anonymous):

Okay thanks